OpenStudy (anonymous):

The square of a certain negative number is equal to five more than one-half of that number. What is the number?

OpenStudy (anonymous):

-2

OpenStudy (asadkarim7):

x is the number square of -x x^2= 1/2 x +5 2x^2 = x + 10 2x^2 -x -10 =0 2x^2 - 5x + 4x-10 = 0 take common x ( 2x -5 ) + 2 ( 2x -5 ) = 0 (x-2)= 0 ; ( 2x-5) =0 x = 2 ; 5/2